Z A 1/3 X 1/2 1/2 у 1/2 х

  1. In the unit cells above, determine the directional indices of the vector, and the Miller Indices of the plane (4 marks)
  1. If the atomic packing factor of a material is known to be 0.5 and there are 5 atoms per unit cell (atomic radius 0.13nm), what is the lattice parameter of the cubic unit cell (3 marks)
  1. Which two of the following atoms would be most likely to combine and form a covalent bond, and briefly explain why (2 marks)

Chlorine (Cl), Germanium (Ge), Calcium (Ca), Chromium (Cr), Carbon (C), Neon (Ne)
Answer #1

a) For the first case,

(1-0)i +(1-1/3)j + (1-0) k = 1i +(2/3)j + 1 k = (1,2/3,1)

directional indices of the vector will be smallest integer representing this = (multiply by 3) = [ 3 2 3]

For the second case, intercept = (1/2, 1/2, 1/2)

Miller indices of the plane = (2,2,2)

a 3 e) u Griven, Atomic No. of atoms, , (h) 5.0 Ra packing, factor? 0.5 0.13 nm 0.13 x 10 -8 con. V APF e so a 2 lattice ve